# Neural Conditional Probability for Inference

### v1

@@ B.3 Proof of Theorem 2
We first bound the error of the conditional expectation model as $\lVert{\mathsf{D}_{Y|X}}-{\widehat{\mathsf{D}}^{\theta}_{Y|X}}\rVert$ as follows
@@
Let Assumption 1 be satisfied. Then there exists an absolute constant $C>0$ such that, for any $\delta\in(0,1)$, it holds w.p.a.l. $1-\delta$
$\lVert\widehat{\mathsf{E}}_{x}[u^{\theta}]-\mathbb{E}[u^{\theta}(X)]\rVert\leq C\left(c_{u}\frac{d\,\log(e\delta^{-1})}{n}+\sigma_{\theta}(X)\sqrt{\frac{\,\log(e\delta^{-1})}{n}}\right).$
Similarly, w.p.a.l. $1-\delta$
$\lVert\widehat{\mathsf{E}}_{y}[v^{\theta}]-\mathbb{E}[v^{\theta}(Y)]\rVert\leq C\left(c_{v}\frac{d\,\log(e\delta^{-1})}{n}+\sigma_{\theta}(Y)\sqrt{\frac{\,\log(e\delta^{-1})}{n}}\right).$
@@
We note that $\lVert Z_{i}\rVert\leq 2c_{u}\,d=:U$ and $\mathrm{Var}(Z_{i})=\mathrm{Var}(\lVert u^{\theta}(X_{i})-\mathbb{E}[u^{\theta}(X_{i})]\rVert)=\sigma^{2}_{\theta}(X)$ for any $i\in[n]$. We apply Minsker, (2017, Corollary 4.1) to get for any $t\geq\frac{1}{6}(U+\sqrt{U^{2}+36n\sigma^{2}_{\theta}(X)})$,
@@
$\lVert\frac{1}{n}\sum_{i=1}^{n}Z_{i}\rVert\leq C\left(\frac{c_{u}\,d}{n}\vee\frac{c_{u}^{2}\,d^{2}}{n^{2}}\vee\frac{\sigma^{3/2}_{\theta}(X)}{n^{3/4}}\vee\frac{\sigma^{2}_{\theta}(X)}{n}+c_{u}\frac{d\,t}{n}+\sigma_{\theta}(X)\sqrt{\frac{t}{n}}\right).$
@@
$\lVert\frac{1}{n}\sum_{i=1}^{n}Z_{i}\rVert\leq C\left(c_{u}\frac{d\,t}{n}\vee\sigma_{\theta}(X)\sqrt{\frac{t}{n}}\right),$
@@ B.4 Sub-Gaussian case
B.5 Estimation of conditional expectation and Conditional covariance
We now derive guarantees for the estimation of the conditional expectation and the conditional covariance for vector-valued output $Y\in\mathbb{R}^{d_{y}}$. We start with a general result for arbitrary vector-valued functions of $Y$. We consider a vector-valued function $\underline{h}=(h_{1},\ldots,h_{\underline{d}})$ where $h_{j}\in L^{2}_{\nu}(\mathcal{Y})$ for any $j\in[\underline{d}]$. We introduce the space of square integrable vector-valued functions $[L^{2}_{\nu}(\mathcal{Y},\mathbb{R}^{\underline{d}})]$ equipped with the norm
@@ B.5 Estimation of conditional expectation and Conditional covariance
$\psi_{n}^{(2)}(\delta)$ $=\epsilon_{n}^{(2)}(\delta)+\left[\psi_{n}(\delta/4)+2(1+\psi_{n}(\delta/4))\varphi_{X}(A)\overline{\epsilon}_{n}(\delta/4)\right]\frac{\mathbb{E}[\lVert Y\rVert^{2}])^{2}}{\sqrt{\mathbb{P}[X\in A]}}$
@@
$\hskip 56.9055pt\leq\left[\psi_{n}(\delta)+2(1+\psi_{n}(\delta))\varphi_{X}(A)\overline{\epsilon}_{n}(\delta)\right]\frac{\mathbb{E}[\lVert Y\rVert^{2}])^{2}}{\sqrt{\mathbb{P}[X\in A]}}.$ (41)
@@ Appendix C Numerical Experiments
Experiments were conducted on a high-performance computing cluster equipped with an Intel(R) Xeon(R) Silver 4210 CPU @ 2.20GHz Sky Lake CPU, 377GB RAM, and an NVIDIA Tesla V100 16Gb GPU.
@@ C.1 Conditional Density Estimation
Method Summary Main hyperparams
Normalizing Flows (NF) (Rezende and Mohamed, 2015b, ) Generative models that transform a simple distribution into a complex one through a series of invertible and differentiable transformations   Architecture   Flow type
Denoising Diffusion Probabilistic Model (DDPM) (Ho et al.,, 2020) Generative models that learn to generate data by reversing a gradual noising process, modeling distributions through iterative refinement   Number of diffusion steps   Noise schedule
Conditional KDE (CKDE) (Li and Racine,, 2006) Nonparametric approach modeling the joint and marginal probabilities via KDE and computes the conditional density as $p(y|x)=p(x,y)/p(x)$.   KDE bandwidth
Mixture Density Network (MDN) (Bishop,, 1994) Uses NeuralNets which takes conditional $x$ as input and governs all the weights of a GMM modeling $p(y|x)$.   NeuralNet architecture   Number of kernels
Kernel Mixture Network (KMN) (Ambrogioni et al.,, 2017) Similar to MDN with the difference that NN only controls the weights of the GMM.   NeuralNet architecture   Method for finding kernel centers   Number of kernels
Least-Squares CDE (LSCDE) (Sugiyama et al.,, 2010) Computes the conditional density as linear combination of Gaussian kernels   Method for finding kernel centers   Number of kernels   Kernels’ bandwidth
Nearest Neighbor Kernel CDE (NNKDE) (Izbicki et al.,, 2017) (Freeman et al.,, 2017) Uses nearest neighbors of the evaluation point $x$ to compute a KDE estimation of $y$.   Number of neighbors   Kernel bandwidth
Random Forest CDE (RFCDE) (Pospisil and Lee,, 2018) Uses a random forest to partition the feature space and constructs a weighted KDE of the output space, based on the weights of the leaves in the forest.   Random forest hyperparams   Basis system   Number of basis
Flexible CDE (FC) (Izbicki and Lee,, 2017) Nonparametric approach which uses a basis expansion of univariate $y$ to turn CDE into a series of univariate regression problems.   Number of expansion coeffs   Regression method hyperparams.
LinCDE (LCDE) (Gao and Hastie,, 2022) Conditional training of unconditional machine learning models to learn density   Number of LinCDE trees

### v2

@@ B.4 Proof of Theorem 2
We first bound the error of the conditional expectation model as $\lVert{\mathsf{D}_{Y|X}}-{\widehat{\mathsf{D}}^{\theta}_{Y|X}}\rVert$ as follows.
@@
Let Assumption 1 be satisfied. Assume in addition that $n\geq c_{u}^{2}d$. Then there exists an absolute constant $C>0$ such that, for any $\delta\in(0,1)$, it holds w.p.a.l. $1-\delta$
$\lVert\widehat{\mathsf{E}}_{x}[u^{\theta}]-\mathbb{E}[u^{\theta}(X)]\rVert\leq C\,c_{u}\sqrt{d}\,\left(\frac{\log(e\delta^{-1})}{n}+\sqrt{\frac{\,\log(e\delta^{-1})}{n}}\right).$
Similarly, if $n\geq c_{v}^{2}d$, w.p.a.l. $1-\delta$
$\lVert\widehat{\mathsf{E}}_{y}[v^{\theta}]-\mathbb{E}[v^{\theta}(Y)]\rVert\leq Cc_{v}\sqrt{d}\left(\frac{\log(e\delta^{-1})}{n}+\sqrt{\frac{\,\log(e\delta^{-1})}{n}}\right).$
@@
We note that $\lVert Z_{i}\rVert\leq 2c_{u}\,\sqrt{d}=:U$ and $\mathrm{Var}(Z_{i})=\mathrm{Var}(\lVert u^{\theta}(X_{i})-\mathbb{E}[u^{\theta}(X_{i})]\rVert)=\sigma^{2}_{\theta}(X)$ for any $i\in[n]$. We apply Minsker, (2017, Corollary 4.1) to get for any $t\geq\frac{1}{6}(U+\sqrt{U^{2}+36n\sigma^{2}_{\theta}(X)})$,
@@
$\lVert\frac{1}{n}\sum_{i=1}^{n}Z_{i}\rVert\leq C\left(\frac{c_{u}\,\sqrt{d}}{n}\vee\frac{c_{u}^{2}\,d}{n^{2}}\vee\frac{\sigma^{3/2}_{\theta}(X)}{n^{3/4}}\vee\frac{\sigma^{2}_{\theta}(X)}{n}+c_{u}\frac{\sqrt{d}\,t}{n}+\sigma_{\theta}(X)\sqrt{\frac{t}{n}}\right).$
@@
$\lVert\frac{1}{n}\sum_{i=1}^{n}Z_{i}\rVert\leq C\,c_{u}\sqrt{d}\left(\frac{t}{n}\vee\sqrt{\frac{t}{n}}\right),$
@@ B.5 Sub-Gaussian case
B.6 Estimation of conditional expectation and Conditional covariance
We now derive guarantees for the estimation of the conditional expectation and the conditional covariance for vector-valued output $Y\in\mathbb{R}^{d_{y}}$.
We start with a general result for arbitrary vector-valued functions of $Y$. We consider a vector-valued function $\underline{h}=(h_{1},\ldots,h_{\underline{d}})$ where $h_{j}\in L^{2}_{\nu}(\mathcal{Y})$ for any $j\in[\underline{d}]$. We introduce the space of square integrable vector-valued functions $[L^{2}_{\nu}(\mathcal{Y},\mathbb{R}^{\underline{d}})]$ equipped with the norm
@@ B.6 Estimation of conditional expectation and Conditional covariance
$\psi_{n}^{(2)}(\delta)$ $=\epsilon_{n}^{(2)}(\delta)+\left[\psi_{n}(\delta/4)+2(1+\psi_{n}(\delta/4))\varphi_{X}(A)\overline{\epsilon}_{n}(\delta/4)\right]\frac{(\mathbb{E}[\lVert Y\rVert^{2}])^{2}}{\sqrt{\mathbb{P}[X\in A]}}$
@@
$\hskip 56.9055pt\leq\left[\psi_{n}(\delta)+2(1+\psi_{n}(\delta))\varphi_{X}(A)\overline{\epsilon}_{n}(\delta)\right]\frac{(\mathbb{E}[\lVert Y\rVert^{2}])^{2}}{\sqrt{\mathbb{P}[X\in A]}}.$ (42)
@@ Appendix C Numerical Experiments
Experiments were conducted on a high-performance computing cluster equipped with an Intel(R) Xeon(R) Silver 4210 CPU @ 2.20GHz Sky Lake CPU, 377GB RAM, and an NVIDIA Tesla V100 16Gb GPU. Code is available at https://github.com/CSML-IIT-UCL/NCP.
@@ C.1 Conditional Density Estimation
Method Summary Main hyperparams Normalizing Flows (NF) (Rezende and Mohamed, 2015b, ) Generative models that transform a simple distribution into a complex one through a series of invertible and differentiable transformations   Architecture   Flow type Denoising Diffusion Probabilistic Model (DDPM) (Ho et al.,, 2020) Generative models that learn to generate data by reversing a gradual noising process, modeling distributions through iterative refinement   Number of diffusion steps   Noise schedule Conditional KDE (CKDE) (Li and Racine,, 2006) Nonparametric approach modeling the joint and marginal probabilities via KDE and computes the conditional density as $p(y|x)=p(x,y)/p(x)$.   KDE bandwidth Mixture Density Network (MDN) (Bishop,, 1994) Uses NeuralNets which takes conditional $x$ as input and governs all the weights of a GMM modeling $p(y|x)$.   NeuralNet architecture   Number of kernels Kernel Mixture Network (KMN) (Ambrogioni et al.,, 2017) Similar to MDN with the difference that NN only controls the weights of the GMM.   NeuralNet architecture   Method for finding kernel centers   Number of kernels Least-Squares CDE (LSCDE) (Sugiyama et al.,, 2010) Computes the conditional density as linear combination of Gaussian kernels   Method for finding kernel centers   Number of kernels   Kernels’ bandwidth Nearest Neighbor Kernel CDE (NNKDE) (Izbicki et al.,, 2017) (Freeman et al.,, 2017) Uses nearest neighbors of the evaluation point $x$ to compute a KDE estimation of $y$.   Number of neighbors   Kernel bandwidth Random Forest CDE (RFCDE) (Pospisil and Lee,, 2018) Uses a random forest to partition the feature space and constructs a weighted KDE of the output space, based on the weights of the leaves in the forest.   Random forest hyperparams   Basis system   Number of basis Flexible CDE (FC) (Izbicki and Lee,, 2017) Nonparametric approach which uses a basis expansion of univariate $y$ to turn CDE into a series of univariate regression problems.   Number of expansion coeffs   Regression method hyperparams. LinCDE (LCDE) (Gao and Hastie,, 2022) Conditional training of unconditional machine learning models to learn density   Number of LinCDE trees
@@ C.2 Confidence Regions
C.3 High-dimensional Experiments
Experiment on high-dimensional synthetic data.
In Fig. 3, we trained NCP for $d=100$ using the same MLP architecture and the same NF with autoregressive flow as in our initial experiments based on $n=10^{5}$ samples $\{(X_{i},Y_{i})\}_{i=1}^{n}$ with values in $\mathbb{R}^{d}\times\mathcal{Y}$. We plot the conditional CDF for several conditioning w.r.t. $\theta(x)$ on 10 repetitions. NCP paired with a small MLP architecture performs comparably to the NF model for Gaussian distributions. For discrete distributions, the NCP demonstrates superior performance compared to the NF model.
We repeated the experiment in Fig. 3 for $d\in\{100,200,500,1000\}$ and recorded the average Kolmogorov-Smirnov (KS) distance of the NCP conditional distribution to the truth, computation time and their standard deviations over 10 repetitions.
100 500 1000
$\theta$ mean std mean std mean std
1.0 0.057 0.019 0.079 0.050 0.062 0.016
1.57 0.041 0.009 0.069 0.042 0.049 0.017
3.14 0.030 0.016 0.116 0.173 0.036 0.010
5.0 0.067 0.052 0.131 0.175 0.072 0.036
Figure 7: Left: we observe only $\approx$ 20% increase in compute time going from $d=10^{2}$ to $d=10^{3}$. Right: average KS distance to the truth and standard deviation over 10 repetitions.
High-dimensional experiment in molecular dynamics: Chignolin folding.
We investigated the dynamics of Chignolin folding, using a molecular dynamics simulation lasting $106\mu s$ and sampled every $200ps$, resulting in $524,743$ data points. Our analysis focuses on 39 heavy atoms (nodes) with a cutoff radius of 5 Angstroms. To predict the conditional transition probability between metastable states, we integrate our NCP approach with a graph neural network (GNN) model. GNNs, as demonstrated by Chanussot et al., (2021), represent the state-of-the-art in modeling atomistic systems, adeptly incorporating the roto-translational and permutational symmetries inherent in physical systems. In particular, we employed a SchNet model Schütt et al., (2019, 2023) with three interaction blocks. Each block features a 64-dimensional latent atomic environment, and the inter-atomic distances for message passing are expanded over $20$ radial basis functions. After the final interaction block, each latent atomic environment is processed through a linear layer and then aggregated by averaging. The model underwent training for $100$ epochs using an Adam optimizer with a learning rate of $10^{-3}$. We employed a batch size of $256$ and set $\gamma$ to $10^{-3}$. In Fig. 2, we show how our NCP approach enables the tracking transitions between metastable states, demonstrating accurate forecasting and strong uncertainty quantification.